OpenStudy (anonymous):

Find all the real zeros of the polynomial. Use the quadratic formula if necessary. If a zero has multiplicity greater than one, only enter the root once. P(x) = x^4 − 8x^3 + 8x^2 + 23x + 6

OpenStudy (anonymous):

ew >.> quartic polynomial. Might as well try to guess a root. Use the Rational Roots Theorem to guess one.

OpenStudy (anonymous):

Isn't the highest exponent the number of zeros?

OpenStudy (anonymous):

Thats right-ish. The highest power tells you how many zeros there are. It is possible that all the zeros could be the same number though. For example: \[x^4+4x^3+6x^2+4x+1 = 0\] Has 4 zeros, they just all happen to be -1. When you factor that polynomial it becomes: \[(x+1)^4 = 0\]

OpenStudy (anonymous):

a huge hint is that since some math teacher made this up, either 1 is a zero or -1 is

OpenStudy (anonymous):

LOLOL

OpenStudy (anonymous):

1 is easy to check, just add the coeffienients: \[1-8+8+23+6\] nope

OpenStudy (anonymous):

i kid you not. it is always either 1 or -1. in this case it is -1

OpenStudy (anonymous):

Best advice ive ever heard on this website lol

OpenStudy (anonymous):

Hey Satellite, where have you been? Some people were ganging up on me yesterday, i needed someone to help me defend my answer >.>

OpenStudy (anonymous):

vacation

OpenStudy (anonymous):

what was it?

OpenStudy (anonymous):

ah, i hope you had a good vacation! It was nothing remotely hard. It was just a bunch of misinformed people. Question was, how many solutions are there to: \[\sqrt{x} = -4\]

OpenStudy (anonymous):

none nil zippo. unless you are working over C

OpenStudy (anonymous):

right right. They were saying x = 16 was a solution, and i was trying my hardest to give every mathematical reason why it wasnt lol.

OpenStudy (anonymous):

what??????

OpenStudy (anonymous):

But no one was buying it. It was like I was some crazy guy trying to spread misinformation and lies >.>

OpenStudy (anonymous):

reason? how about the fact that \[\sqrt{16}=4\] and last time i checked \[4\neq -4\]

OpenStudy (anonymous):

I told them how if you square equations you create what are called extraneous solutions, and that x = 16 was extraneous. I told them that the graphs of f(x) = sqrtx and g(x) = -4 dont intersect. They didnt listen to any of it lol.

OpenStudy (anonymous):

oh i know this one. it goes like this: \[(-4)^2=16\] so \[\sqrt{16}=\pm4\] i have battled against this before

OpenStudy (anonymous):

Yeah, they kept bring that up, and i was like, "yes, i know that, but square roots only return positive values. If you want the negative, you have to denote that with a negative sign. Since this problem doesnt have a negative sign, its not asking for it." They proceeded to mock me anyways!

OpenStudy (anonymous):

None of my friends were around to help (or anyone willing to listen to reasoning, or anyone semi-intelligent for that matter <.<) :(

OpenStudy (anonymous):

well let them suffer under their false delusions. apparently the square root function for them is not well defined. i have had to resort to "you can bank on it" for this argument so let them mock and be mistaken. now dinner time

OpenStudy (anonymous):

lolol, thanks, i just had to get that off my chest haha